OpenStudy (anonymous):

multiply (b+1/3)(b+1/2)

OpenStudy (anonymous):

= b^2 + 1/3 b + 1/2 b + 1/6 = b^2 5/6 b + 1/6

OpenStudy (anonymous):

sorry thats b^2 + 5/6 b + 1/6

OpenStudy (anonymous):

\[(b+1/3)*(b+1/2)= b ^{2}+(5b/6)+1/6\]